40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 28 20:42:53 2020, Koji, Update, General, Vent Prep 10x
    Reply  Thu Oct 29 12:43:35 2020, gautam, Update, General, Vent Prep preVentAlignment.pngPRMIcarr_PRMalignment.pngALS_ool.pdflowPowerMC.png
       Reply  Thu Oct 29 19:30:56 2020, gautam, Update, General, Vent complete ventValveStatus.pngvent_Oct2020.jpgventValveStatus_overnight.pnglowPowerIMC_afterVent.png
          Reply  Mon Nov 2 12:44:25 2020, gautam, Update, General, ETMY and OMC chamber heavy doors off CB84FD62-C48B-414F-8825-50C9DE0B47CF.jpeg
             Reply  Mon Nov 2 16:46:06 2020, gautam, Update, General, Plan for OMC chamber IMG_2318.JPGIMG_2332.JPG
                Reply  Mon Nov 2 17:13:19 2020, Koji, Update, General, Plan for OMC chamber 
                   Reply  Mon Nov 2 17:32:05 2020, gautam, Update, General, Plan for OMC chamber IMG_2317.JPG
Message ID: 15649     Entry time: Wed Oct 28 20:42:53 2020     Reply to this: 15651
Author: Koji 
Type: Update 
Category: General 
Subject: Vent Prep 
  1. Two arms / BS / PRM / SRM were aligned. (Attachment 1)
  2. IMC was aligned by WFS and the WFS offsets were offloaded.
  3. Suspension Status Snapshot (Attachment 2)
  4. Oplevs are aligned (Attachment 3)
  5. Xarm green was aligned in the daytime. Xarm green refl DC (C1:ALS-X_REFL_DC_OUTPUT) was 620 (aligned) ~1300 (drifted). When unlocked, it was 3750.
  6. Yarm green: I saw no flash. We don't have functional PZT alignment since the ASY M2 PZT got broken. I went to the Yend. Something funky is going on with the Yend green. I struggled to have any flash of the cavity. The apertures were not so precise. I finally got TEM00 locked, but the modematching seems exteremely low (like 1/1000?). Basically I saw no power reduction of the refl when the cavity is locked. So at least the cavity was locked but we might need to revisit when we open the chamber
    ==> Gautam thinks it was not like that. So he will check the green alignment tomorrow (Thu).
  7. Item checking: I familialized myself with the yend crane operation. Today I learned that there is a power switch on the wall (Attachment 4). The yend has two heavy door storages (Attachments 4/5). The slings to lift the heavy door are in the crane cabinet along with the y arm (Attachment 6). I didn't yet try to find the "hammer" to hit the door if the o-ring stuck too strong, although that's optional.
  8. We want to reduce the PSL power. But Gautam wants to use the arm locking with the nominal power, it will be done tomorrow by him.
     
  9. The last thing is to check the green trans power. I noticed that the green trans beams are blocked by an HWP for the BHD LO path on the PSL table. (Attachment 7)
    The HWP was moved and the process was recorded in the movie (Attachment 8). The fiber output was monitored by the BHD DC (aka AS110 DC) with the AS path blocked. The fiber output of 22.6mV (offset -2.5mV) was improved tio 29.1mV after the HWP move and the alignment adjustment.
  10. Now the green transmissions are visible by the green PDs. Attachment 9 shows the trans and ref of each green beams with and without locking to TEM00. The questionable green TRY was ~0.3. If we compare this with the histrical data (Attachment 10), it is about 1/4 of the value in the past. It's not too crazy but still quite low.

At this point, I'm leaving the lab. All the suspensions (incl SRM) are aligned. PSL/GRX/GRY shutters were left open.

Attachment 1: Screen_Shot_2020-10-28_at_19.42.27.png  193 kB  Uploaded Wed Oct 28 21:44:14 2020  | Hide | Hide all
Screen_Shot_2020-10-28_at_19.42.27.png
Attachment 2: Screen_Shot_2020-10-28_at_19.42.38.png  301 kB  Uploaded Wed Oct 28 21:44:31 2020  | Hide | Hide all
Screen_Shot_2020-10-28_at_19.42.38.png
Attachment 3: Screen_Shot_2020-10-28_at_20.42.26.png  296 kB  Uploaded Wed Oct 28 21:44:34 2020  | Hide | Hide all
Screen_Shot_2020-10-28_at_20.42.26.png
Attachment 4: P_20201028_222839.jpg  620 kB  Uploaded Thu Oct 29 00:50:05 2020  | Hide | Hide all
P_20201028_222839.jpg
Attachment 5: P_20201028_223020.jpg  604 kB  Uploaded Thu Oct 29 00:50:09 2020  | Hide | Hide all
P_20201028_223020.jpg
Attachment 6: P_20201028_223212.jpg  627 kB  Uploaded Thu Oct 29 00:50:13 2020  | Hide | Hide all
P_20201028_223212.jpg
Attachment 7: 2020-10-28-233038.jpeg  1.777 MB  Uploaded Thu Oct 29 00:58:02 2020  | Hide | Hide all
2020-10-28-233038.jpeg
Attachment 8: HWP_moving.mp4  30.365 MB  Uploaded Thu Oct 29 03:09:25 2020
Attachment 9: Screenshot_from_2020-10-29_01-04-06.png  83 kB  Uploaded Thu Oct 29 03:09:44 2020  | Hide | Hide all
Screenshot_from_2020-10-29_01-04-06.png
Attachment 10: Screenshot_from_2020-10-29_01-16-53.png  186 kB  Uploaded Thu Oct 29 03:09:48 2020  | Hide | Hide all
Screenshot_from_2020-10-29_01-16-53.png
ELOG V3.1.3-